# Hybrid Quantum Repeater Chains with Semiconductor Quantum Dots and Group-IV-Vacancy Color Centers in Diamond

### v2

@@ Abstract
Figure 1: QD-G4V hybrid repeater chain protocol. a) Elementary-link establishment. The chain consists of two end nodes, Alice and Bob, and N1N-1 intermediate repeater nodes, each represented by an orange rectangle color block. QD emitters (red stars) are positioned midway between neighboring nodes and distribute entangled photons (red bell shapes) between them. In each node, G4V electron spins (dark blue) serve as interfaces for loading qubits from the incoming entangled photons, while nuclear spins (orange) provide long-term qubit storage. b) The fundamental operations of the elementary-link entanglement distillation. They involve local nucleus-nucleus cnot gates and qubit measurement and classical communications between the two nodes. c) The entanglement distillation protocol following Ref. [1]. It has four options: level-0: using directly the raw Bell pairs; level-1, -2, and -3: each distilled Bell pair is made up from 2, 4, and 6 raw ones, respectively. Single-qubit gates (not shown) are applied before the level-2 and -3 distillations. d) Entanglement swapping. It contains a cnot gate between the two nuclear spins from different memory modules in a repeater node, followed by a Hadamard gate and qubit measurements. e) The end-to-end entanglement distillation is similar to the elementary-link distillation, except that, due to its time-consuming nature, classical communication is employed for post-selection rather than as a precondition in the level-2 and level-3 operations.
@@
Figure 2: Optical circuits for the repeater-chain operations. a) Entanglement transfer from the photons to the nuclear spins as in the elementary-link establishment step. QDs emit frequency-distinct, polarization-entangled photon pairs, which are separated by a wavelength-division multiplexer (WDMP). Frequency filters ($F$s) are used on one side or both for narrowing the bandwidth of the photon(s) in a Bell pair. Polarization-to-time-bin converters (PTBCs), implemented with polarizing beam splitters, encode polarization qubits into time-bin qubits for fiber transmission. At each repeater node, frequency converters (FCs) shift photons from telecommunication to visible wavelengths for G4V compatibility. Optical switches (S) direct photons to designated memory cells. Through controlled photon–spin interactions and subsequent photon detection, the photon qubit is mapped onto a G4V electron spin (dark blue) and then transferred to a nuclear spin (orange). Successful detection heralds photon receipt, which is confirmed via classical communication between nodes. b) and c) Optical circuit for the local nucleus-nucleus cont gates in the entanglement distillation and swapping steps, respectively. A single photon source (implemented with a separate G4V [2]) generates a photon, which is guided by circulators and optical switches (S). It interacts sequentially with two target G4V electron spins until heralded photon measurements confirm entanglement between them. Nucleus–nucleus cnot gates are then realized using local electron–nuclear gates.
@@ I Introduction
Quantum networks [3, 4, 5, 6, 7, 8] capable of distributing high-fidelity entanglement are essential for applications such as distributed quantum sensing [9, 10, 11, 12], secure communication [13, 14, 15], and networking of quantum computers [16, 17]. A major challenge for the construction of extended quantum networks is photon loss in optical fibers: 9999% of photons are lost over every $100$ km of transmission [18]. Since quantum states cannot be amplified like classical signals [19], quantum repeaters [20] are required to extend entanglement over continental distances.
A variety of hardware platforms have been explored for quantum repeaters, including atomic ensembles [21], trapped ions [22], solid-state defect centers [23], and semiconductor quantum dots [24]. Each offers distinct strengths: quantum dots (QDs) can generate entangled photons at high rates with excellent fidelity [25], while group-IV-vacancy (G4V) centers in diamond provide long-lived, optically addressable spins for efficient and heralded photon storage [26, 27]. These complementary capabilities naturally motivate hybrid repeater architectures that combine the best features of different hardware systems.
Here, we propose and analyze such a hybrid scheme, integrating QDs as entangled photon-pair sources with G4V centers as efficient single-spin memories. A key challenge is the bandwidth mismatch: QDs emit photons with $\sim$10 GHz bandwidths [28, 29, 25], while G4Vs have narrow optical transitions of order 100 MHz [30, 31]. To address this challenge, we use a comprehensive model in [32] of the cavity-mediated spin–photon interaction that includes polarization effects, magnetic-field-induced level splitting, and cross-talk between transitions. Having expanded beyond more simplified models [33], our approach enables the identification of practical mitigation strategies and provides a realistic assessment of their effectiveness. In particular, we investigate techniques such as frequency filtering and magnetic-field optimization to determine the conditions under which efficient, high-fidelity photon storage can be achieved.
Building on this, we evaluate the performance of a full repeater chain with entanglement distillation using an operational protocol based on BB84 quantum key distribution [34]. Our results show that a network with thousands of G4V memories across multiple nodes can reach secret-key rates on the order of hundreds to thousands of bits per second over continental distances. This demonstrates the strong potential of this hybrid QD–G4V platform as a path toward scalable long-distance quantum networks.
@@ II Hardware Implementation
The hybrid repeater chain protocol is illustrated in Fig. 1. The two end nodes, Alice and Bob, are separated by a large distance. The repeater nodes divide the total distance into multiple segments, each containing a QD emitter in the middle for sending photons to the two neighboring nodes. Each repeater node contains two memory modules, one of each segment. Depending on whether entanglement distillation is applied, the end node contains one or no memory modules. When there is no entanglement distillation, the received photons are directly measured, hence no quantum memories at the end nodes are required. Each memory module contains multiple memory cells, each of which contains a G4V electron spin serving as the communication qubit and a nuclear spin, e.g., 13[35, 36], 29Si [37], 117Sn [38], or 119Sn, serving as the storage qubit. Here, the G4V electron spin refers to the collection of electrons in the unsaturated covalent bonds [39, 40]. It is effectively equivalent to a hole with spin-$\tfrac{1}{2}$.
Our two-way repeater chain protocol consists of three key operations: elementary link establishment, entanglement distillation [41, 42, 43, 44, 42, 1], and entanglement swapping [45, 21]. Elementary link establishment generates entanglement between neighboring nodes separated by a short distance. Entanglement distillation enhances the fidelity of Bell pairs at the cost of consuming multiple pairs. Two possibilities, entanglement distillation for the elementary links and the end-to-end links, are examined in this work. Entanglement swapping extends entanglement by merging two adjacent links into a longer one. We describe these three operations in detail below.
@@ II.1 Elementary-Link Establishment
The quantum dot (QD) generates entangled photon pairs with polarization correlations [46, 47, 48, 49, 25, 50]. Photon absorption and emission in the QD correspond to an electron transitioning between the conduction and valence bands. As a result, the four lowest-energy quantum states of the QD are: the ground state $|0\rangle$, the two single-exciton states $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$ of opposite spin orientations, and the biexciton state $|\rm XX\rangle$, as illustrated in Fig. 3a. Under suitable laser irradiation, the QD is initialized in the biexciton state $|\rm XX\rangle$. From this state, it can decay via two possible cascaded paths: $|\rm XX\rangle\rightarrow|\rm X\rangle\rightarrow|0\rangle$ and $|\rm XX\rangle\rightarrow|\rm X^{\prime}\rangle\rightarrow|0\rangle$. The first (second) path produces two vertically (horizontally) polarized photons of different frequencies.
@@
The two emitted photons have different frequencies and can thus be separated using a wavelength-division multiplexer (WDMP) [51] and be directed to the two different nodes of the segment. To mitigate polarization instability during fiber transmission, before transmission, a polarization-to-time-bin converter (PTBC) [52] is used to convert the polarization states $|\rm H\rangle$ and $|\rm V\rangle$ ($|\rm H^{\prime}\rangle$ and $|\rm V^{\prime}\rangle$) into early and late time-bin states $|\rm E\rangle$ and $|\rm L\rangle$ ($|\rm E^{\prime}\rangle$ and $|\rm L^{\prime}\rangle$), respectively.
After the polarization-to-time-bin conversion, the two photons will be sent to the two nodes adjacent to the QD emitter, respectively. When a photon arrives at a repeater node, a frequency converter [53] shifts its frequency from the telecom band to the frequency of the G4V spin-photon interface in the visible light regime. The photon is then routed to a designated memory cell by an optical switching network. Through a tailored interaction between the photon and the target G4V electron spin [54, 8], the quantum information (qubit) carried by the photon is transferred to the electron spin. This interaction is discussed in detail further down.
Photons experience significant loss during transmission through optical fibers and on-chip circuits. Therefore, heralding their successful arrival is essential for identifying the memory cells that have been successfully entangled and are ready for the subsequent operation of entanglement swapping. At each repeater node, local heralding is achieved by measuring the photon in the $X$-basis, $|\pm\rangle=\frac{1}{\sqrt{2}}(|\rm E\rangle\pm|\rm L\rangle)$, after the quantum state transfer operation [54, 8], i.e. the interaction between the photon and the target electron spin. Moreover, a successfully transferred qubit is further transferred to a nuclear spin via an electron-nuclear entanglement gate [36] for long-time storage. Lastly, to determine whether the other photon of the same pair has also been received at the opposite node, classical communication between the two nodes is required to exchange heralding signals.
@@ II.2 Entanglement distillation and swapping
We examine the possibilities to apply entanglement distillations for the elementary entanglement links and the end-to-end links as shown in Fig. 1b and e. The distillation protocol we adopted is presented in Ref. [1]. It introduces three levels of purification. By applying single-qubit rotations, different error types in the entangled state are corrected at successive levels. A level-1, -2, and -3 distilled Bell pair requires at least two, four, and six raw Bell pairs, respectively, as shown in Fig. 1c. The basic operation in this scheme is the fusion of two Bell pairs into one. Firstly, two pairs of remotely entangled qubits are established between two nodes. Then a cnot gate is implemented locally at each node.
A direct nuclear-nuclear cnot gate could be implemented if both spins belong to the same crystal and are coupled to a comon G4V center [55]. However, due to extensive optical multiplexing requirement, we consider the local nuclear-nuclear cnot gates mediated by the two G4V electron spins and additional photons. The reason is as follows. While coupling to tens of 13C spins has been demonstrated in NV centers [35], high-fidelity operations are typically feasible only with a small subset of strongly coupled spins. Moreover, because G4V centers exhibit long reset times and slower G4V–13C gates compared to fast quantum dot emitters, hundreds of G4Vs are required for sufficient multiplexing. Thus, the probability that two targeted spins reside at the same site is low, making inter-G4V operations necessary for scalable architectures. We therefore focus on on a scenario where each nuclear spin is hosted by a separate G4V.
A schematic optical circuit for the fusion operation is shown in Fig. 2b. The procedure consists of two steps. First, entanglement is generated between the two electron spins [56]. This step can be retried until either success or a maximum number of attempts is reached. A cap in the number of attempts ensures synchronization across the repeater chain and limits decoherence in the nuclear memory. In each attempt, a single photon—emitted by a G4V source or another compatible emitter—is prepared as a time-bin qubit and routed through an optical circuit containing two circulators. The photon sequentially interacts with the electron spins of both G4V-memories before being measured in the $X$ basis. Successful detection heralds entanglement between the electron spins. If no heralding event is observed, the G4Vs are re-initialized while the nuclear spin states remain preserved.
Once entanglement between two electron spins is established, the second step performs local Bell-state measurements between each electron and its adjacent nuclear spin. These measurements, enabled by electron–nuclear entangling gates, realize a near-deterministic cnot gate between the two nuclear spins [36].
@@ III System compatibility
Figure 3: QD and G4V hardware profiles. a) Energy levels and decay paths of a quantum dot (QD). The QD includes a ground state $|0\rangle$, two approximately degenerate single-excitation states, $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$, and a biexcitation state, $|\rm XX\rangle$, which together give rise to two cascade decay paths. The decay times $T_{\rm X}$ and $T_{\rm XX}$ of the single- and biexcitation states correspond to linewidths $\gamma_{\rm X}$ and $\gamma_{\rm XX}$, respectively. Photons are emitted with horizontal (H) or vertical (V) polarization, depending on the decay path; the presence or absence of a prime symbol for the labels H and V denotes different photon frequencies. b) Frequency profile of photons emitted by the QD. According to the experiment presented in [25], the photons from both $|\rm XX\rangle$, $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$ exhibit Lorentzian spectral profiles centered at $\omega_{\rm 0,X}$ and $\omega_{\rm 0,XX}$, with bandwidths $\gamma_{\rm XX}=8.33$ GHz and $\gamma_{\rm X}=4.34$ GHz, respectively. c) Fiber–cavity interface and integration of a group-IV-vacancy color center (G4V) into the sawfish nanophotonic crystal cavity [31]. The incoming photon from the fiber is in a superposition of early and late time-bin states, encoding a qubit. The cavity supports a single optical mode with resonance frequency $\omega_{c}$, coupling strength $\bm{g}=[g_{1A},g_{2B},g_{2A},g_{1B}]$ to the G4V, and a total loss rate $\kappa$. d) Energy levels of the G4V center (SiV or SnV), modeled as a four-level system with ground states $|1\rangle$ and $|2\rangle$ and excited states $|A\rangle$ and $|B\rangle$. The optical transitions $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ have linewidths $\gamma_{1A}$ and $\gamma_{2B}$, respectively. The cavity resonance $\omega_{c}$ is detuned by $\delta$ from the $|1\rangle\leftrightarrow|A\rangle$ transition. The ground-states $|1\rangle$ and $|2\rangle$ circled by the orange line make up the qubit of the G4V spin, whose splitting is denoted as $\omega_{s}$. e) Operational sequence for implementing the spin–photon entanglement gate via the reflection-based scheme (see App. A). The protocol consists of four steps: (1) initialize the G4V spin in state $|1\rangle$; (2) scatter the early time-bin photon; (3) apply a $\pi/2$ rotation around the $y$-axis to the spin; (4) scatter the late time-bin photon. Fig. e) is adapted from [54].
@@
Compatibility between the QD photon emitter and the G4V single-spin quantum memory is a central challenge in our protocol. While the central frequencies of the two systems can be aligned using frequency converters, matching their bandwidths and lineshapes remains a formidable task. The QD emitter has a natural lifetime of 125125250250 ps [25], corresponding to a bandwidth on the order of 10 GHz. In contrast, the G4V spin has a natural lifetime of 1.71.74.54.5 ns [31], implying a bandwidth on the order of 100 MHz—only about 1% of that of the QD. This mismatch implies that the G4V would fail to interact efficiently with approximately 99% of the QD-emitted photons. Nevertheless, this issue can be mitigated by coupling the G4V and the broadband photon via an intermediate cavity with an optimized total energy decay rate and central mode frequency relative to the G4V transition [33].
Following a decay cascade from the biexcitation state to the ground state, the experimentally investgated QD considered here emits two photons with Lorentzian spectral profiles [57] and bandwidths $\gamma_{\rm XX}=8.33$ GHz and $\gamma_{\rm X}=4.34$ GHz, respectively [25], as shown in Fig. 3(b). Each photon can optionally pass through a Fabry–Perot interferometer for frequency filtering [58] (see App. D.4 for details) and is routed to one of the two adjacent nodes containing a G4V spin, thereby distributing entanglement. Since the two photons have different bandwidths, using two distinct cavities optimized for each photon would yield the best performance. However, for simplicity and practical implementation, we adopt a symmetric design for both cavities, balancing system complexity and performance. As a result, only three parameters require optimization: the photon central frequency $\omega_{0}$ (as targeted by the frequency converters), the cavity resonance frequency $\omega_{c}$, and the cavity loss rate $\kappa$.
Because the challenge of low success probability of entanglement transfer, mainly caused by photon loss during fiber transmission and local optical circuits, is addressed through multiplexing, quantum state and gate fidelities become the dominant factor limiting repeater performance. Therefore, our objective is to maximize the fidelity of the resulting G4V spin-spin state relative to the raw G4V electron-electron Bell-state, $|\mathrm{Bell}\rangle=\frac{1}{\sqrt{2}}(|11\rangle+|22\rangle)$, where $|ij\rangle$ ($i,j\in\{1,2\}$) is the quantum state where the left(right) G4V electron in an elementary link is in the $|i(j)\rangle$ state. This leads to the objective function [59]
@@
To examine the quantum state transfer from one of the two photons in the pair to the G4V spin, we consider the setup shown in Fig. 3(c). Via the polarization to time-bin converter, the photonic qubit is encoded in early and late time bins. The photon is guided through a fiber into a single-sided, overcoupled nanophotonic crystal cavity [31], where it interacts with the G4V spin. After interaction, the photon is reflected back into the fiber and directed to a measurement device. In the simplest case spin-photon entanglement succeeds when the phase contrast fulfills $\Delta\phi(\omega):=\phi_{1}(\omega)-\phi_{2}(\omega)=\pi$ on a frequency range around the central frequency of the incoming photon where the expressions $\phi_{i}$ are the phases of the reflected photon when the spin is initialized in $|i\rangle$ for $i=1,2$. The phase contrast $\Delta\phi$ is determined by the optical splitting $\Delta\omega_{s}$ of the G4V, which corresponds to the energy difference between the optical transitions connecting the excited states $|A\rangle$ and $|B\rangle$, and the ground states $|1\rangle$ and $|2\rangle$.

### v3

@@ Abstract
Figure 1: QD-G4V hybrid repeater chain protocol. a) Elementary-link establishment. The chain consists of two end nodes, Alice and Bob, and N1N-1 intermediate repeater nodes, each represented by an orange rectangle color block. QD emitters (red stars) are positioned midway between neighboring nodes and distribute entangled photons (red bell shapes) between them. In each node, G4V electron spins (dark blue) serve as interfaces for loading qubits from the incoming entangled photons, while nuclear spins (orange) provide long-term qubit storage. b) The fundamental operations of the elementary-link entanglement distillation. They involve local nucleus-nucleus cnot gates and qubit measurement and classical communications between the two nodes. c) The entanglement distillation protocol following Ref. [distillation_nigmatullin2016minimally]. It has four options: level-0: using directly the raw Bell pairs; level-1, -2, and -3: each distilled Bell pair is made up from 2, 4, and 6 raw ones, respectively. Single-qubit gates (not shown) are applied before the level-2 and -3 distillations. d) Entanglement swapping. It contains a cnot gate between the two nuclear spins from different memory modules in a repeater node, followed by a Hadamard gate and qubit measurements. e) The end-to-end entanglement distillation is similar to the elementary-link distillation, except that, due to its time-consuming nature, classical communication is employed for post-selection rather than as a precondition in the level-2 and level-3 operations.
@@
Figure 2: Optical circuits for the repeater-chain operations. a) Entanglement transfer from the photons to the nuclear spins as in the elementary-link establishment step. QDs emit frequency-distinct, polarization-entangled photon pairs, which are separated by a wavelength-division multiplexer (WDMP). Frequency filters ($F$s) are used on one side or both for narrowing the bandwidth of the photon(s) in a Bell pair. Polarization-to-time-bin converters (PTBCs), implemented with polarizing beam splitters, encode polarization qubits into time-bin qubits for fiber transmission. At each repeater node, frequency converters (FCs) shift photons from telecommunication to visible wavelengths for G4V compatibility. Optical switches (S) direct photons to designated memory cells. Through controlled photon–spin interactions and subsequent photon detection, the photon qubit is mapped onto a G4V electron spin (dark blue) and then transferred to a nuclear spin (orange). Successful detection heralds photon receipt, which is confirmed via classical communication between nodes. b) and c) Optical circuit for the local nucleus-nucleus cont gates in the entanglement distillation and swapping steps, respectively. A single photon source (implemented with a separate G4V [Knall2022]) generates a photon, which is guided by circulators and optical switches (S). It interacts sequentially with two target G4V electron spins until heralded photon measurements confirm entanglement between them. Nucleus–nucleus cnot gates are then realized using local electron–nuclear gates.
@@ I Introduction
Quantum networks [Zhang2024, wei_towards_2022, Sukachev2021, ruf_quantum_2021, lodahl_quantum-dot_2017, reiserer_cavity-based_2015] capable of distributing high-fidelity entanglement are essential for applications such as distributed quantum sensing [kim_distributed_2024, zhang_distributed_2021, guo_distributed_2020, Degen2017], secure communication [Ramya2025, Chen2021, gisin_quantum_2007], and networking of quantum computers [sunami_scalable_2025, aghaee_rad_scaling_2025]. A major challenge for the construction of extended quantum networks is photon loss in optical fibers: 9999% of photons are lost over every $100$ km of transmission [miya_ultimate_1979]. Since quantum states cannot be amplified like classical signals [wootters_single_1982], quantum repeaters [briegel_quantum_1998] are required to extend entanglement over continental distances.
A variety of hardware platforms have been explored for quantum repeaters, including atomic ensembles [Sangouard2011], trapped ions [Krutyanskiy2023], solid-state defect centers [Wang2023], and semiconductor quantum dots [Wang2012]. Each offers distinct strengths: quantum dots (QDs) can generate entangled photons at high rates with excellent fidelity [schimpf_quantum_2021], while group-IV-vacancy (G4V) centers in diamond provide long-lived, optically addressable spins for efficient and heralded photon storage [knaut_quantum_2024, harris_coherence_2023]. These complementary capabilities naturally motivate hybrid repeater architectures that combine the best features of different hardware systems.
Here, we propose and analyze such a hybrid scheme, integrating QDs as entangled photon-pair sources with G4V centers as efficient single-spin memories. A key challenge is the bandwidth mismatch: QDs emit photons with $\sim$10 GHz bandwidths [Neuwirth2021, zajac_quantum_2025, schimpf_quantum_2021], while G4Vs have narrow optical transitions of order 100 MHz [Sukachev2017, bopp_sawfish_2024]. To address this challenge, we use a comprehensive model in [strocka_memory_2025] of the cavity-mediated spin–photon interaction that includes polarization effects, magnetic-field-induced level splitting, and cross-talk between transitions. Having expanded beyond more simplified models [omlor_entanglement_2024], our approach enables the identification of practical mitigation strategies and provides a realistic assessment of their effectiveness. In particular, we investigate techniques such as frequency filtering and magnetic-field optimization to determine the conditions under which efficient, high-fidelity photon storage can be achieved.
Building on this, we evaluate the performance of a full repeater chain with entanglement distillation using an operational protocol based on BB84 quantum key distribution [shor_simple_2000]. Our results show that a network with thousands of G4V memories across multiple nodes can reach secret-key rates on the order of hundreds to thousands of bits per second over continental distances. This demonstrates the strong potential of this hybrid QD–G4V platform as a path toward scalable long-distance quantum networks.
@@ II Hardware Implementation
The hybrid repeater chain protocol is illustrated in Fig. 1. The two end nodes, Alice and Bob, are separated by a large distance. The repeater nodes divide the total distance into multiple segments, each containing a QD emitter in the middle for sending photons to the two neighboring nodes. Each repeater node contains two memory modules, one of each segment. Depending on whether entanglement distillation is applied, the end node contains one or no memory modules. When there is no entanglement distillation, the received photons are directly measured, hence no quantum memories at the end nodes are required. Each memory module contains multiple memory cells, each of which contains a G4V electron spin serving as the communication qubit and a nuclear spin, e.g., 13[Tim2024mapping, beukers_control_2024], 29Si [phone_gate_stas2022robust], 117Sn [Parker2023], or 119Sn, serving as the storage qubit. Here, the G4V electron spin refers to the collection of electrons in the unsaturated covalent bonds [gu2025group, Hepp2014SiVthesis]. It is effectively equivalent to a hole with spin-$\tfrac{1}{2}$.
Our two-way repeater chain protocol consists of three key operations: elementary link establishment, entanglement distillation [rozpedek_optimizing_2018, bratzik_quantum_2013, horodecki_quantum_2009, dur_quantum_1999, bratzik_quantum_2013, distillation_nigmatullin2016minimally], and entanglement swapping [yan_survey_2021, Sangouard2011]. Elementary link establishment generates entanglement between neighboring nodes separated by a short distance. Entanglement distillation enhances the fidelity of Bell pairs at the cost of consuming multiple pairs. Two possibilities, entanglement distillation for the elementary links and the end-to-end links, are examined in this work. Entanglement swapping extends entanglement by merging two adjacent links into a longer one. We describe these three operations in detail below.
@@ II.1 Elementary-Link Establishment
The quantum dot (QD) generates entangled photon pairs with polarization correlations [QD_1_reimer2012bright, QD_2_heinze2015quantum, QD_3_muller2014demand, QD_4_strobel2024high, schimpf_quantum_2021, Huber2018]. Photon absorption and emission in the QD correspond to an electron transitioning between the conduction and valence bands. As a result, the four lowest-energy quantum states of the QD are: the ground state $|0\rangle$, the two single-exciton states $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$ of opposite spin orientations, and the biexciton state $|\rm XX\rangle$, as illustrated in Fig. 3a. Under suitable laser irradiation, the QD is initialized in the biexciton state $|\rm XX\rangle$. From this state, it can decay via two possible cascaded paths: $|\rm XX\rangle\rightarrow|\rm X\rangle\rightarrow|0\rangle$ and $|\rm XX\rangle\rightarrow|\rm X^{\prime}\rangle\rightarrow|0\rangle$. The first (second) path produces two vertically (horizontally) polarized photons of different frequencies.
@@
The two emitted photons have different frequencies and can thus be separated using a wavelength-division multiplexer (WDMP) [aoyama_optical_1979] and be directed to the two different nodes of the segment. To mitigate polarization instability during fiber transmission, before transmission, a polarization-to-time-bin converter (PTBC) [kupchak_time-bin--polarization_2017] is used to convert the polarization states $|\rm H\rangle$ and $|\rm V\rangle$ ($|\rm H^{\prime}\rangle$ and $|\rm V^{\prime}\rangle$) into early and late time-bin states $|\rm E\rangle$ and $|\rm L\rangle$ ($|\rm E^{\prime}\rangle$ and $|\rm L^{\prime}\rangle$), respectively.
After the polarization-to-time-bin conversion, the two photons will be sent to the two nodes adjacent to the QD emitter, respectively. When a photon arrives at a repeater node, a frequency converter [albrecht_waveguide_2014] shifts its frequency from the telecom band to the frequency of the G4V spin-photon interface in the visible light regime. The photon is then routed to a designated memory cell by an optical switching network. Through a tailored interaction between the photon and the target G4V electron spin [bhaskar_experimental_2020, reiserer_cavity-based_2015], the quantum information (qubit) carried by the photon is transferred to the electron spin. This interaction is discussed in detail further down.
Photons experience significant loss during transmission through optical fibers and on-chip circuits. Therefore, heralding their successful arrival is essential for identifying the memory cells that have been successfully entangled and are ready for the subsequent operation of entanglement swapping. At each repeater node, local heralding is achieved by measuring the photon in the $X$-basis, $|\pm\rangle=\frac{1}{\sqrt{2}}(|\rm E\rangle\pm|\rm L\rangle)$, after the quantum state transfer operation [bhaskar_experimental_2020, reiserer_cavity-based_2015], i.e. the interaction between the photon and the target electron spin. Moreover, a successfully transferred qubit is further transferred to a nuclear spin via an electron-nuclear entanglement gate [beukers_control_2024] for long-time storage. Lastly, to determine whether the other photon of the same pair has also been received at the opposite node, classical communication between the two nodes is required to exchange heralding signals.
@@ II.2 Entanglement distillation and swapping
We examine the possibilities to apply entanglement distillations for the elementary entanglement links and the end-to-end links as shown in Fig. 1b and e. The distillation protocol we adopted is presented in Ref. [distillation_nigmatullin2016minimally]. It introduces three levels of purification. By applying single-qubit rotations, different error types in the entangled state are corrected at successive levels. A level-1, -2, and -3 distilled Bell pair requires at least two, four, and six raw Bell pairs, respectively, as shown in Fig. 1c. The basic operation in this scheme is the fusion of two Bell pairs into one. Firstly, two pairs of remotely entangled qubits are established between two nodes. Then a cnot gate is implemented locally at each node.
A direct nuclear-nuclear cnot gate could be implemented if both spins belong to the same crystal and are coupled to a comon G4V center [reyes2022completeBell]. However, due to extensive optical multiplexing requirement, we consider the local nuclear-nuclear cnot gates mediated by the two G4V electron spins and additional photons. The reason is as follows. While coupling to tens of 13C spins has been demonstrated in NV centers [Tim2024mapping], high-fidelity operations are typically feasible only with a small subset of strongly coupled spins. Moreover, because G4V centers exhibit long reset times and slower G4V–13C gates compared to fast quantum dot emitters, hundreds of G4Vs are required for sufficient multiplexing. Thus, the probability that two targeted spins reside at the same site is low, making inter-G4V operations necessary for scalable architectures. We therefore focus on on a scenario where each nuclear spin is hosted by a separate G4V.
A schematic optical circuit for the fusion operation is shown in Fig. 2b. The procedure consists of two steps. First, entanglement is generated between the two electron spins [Knaut2024]. This step can be retried until either success or a maximum number of attempts is reached. A cap in the number of attempts ensures synchronization across the repeater chain and limits decoherence in the nuclear memory. In each attempt, a single photon—emitted by a G4V source or another compatible emitter—is prepared as a time-bin qubit and routed through an optical circuit containing two circulators. The photon sequentially interacts with the electron spins of both G4V-memories before being measured in the $X$ basis. Successful detection heralds entanglement between the electron spins. If no heralding event is observed, the G4Vs are re-initialized while the nuclear spin states remain preserved.
Once entanglement between two electron spins is established, the second step performs local Bell-state measurements between each electron and its adjacent nuclear spin. These measurements, enabled by electron–nuclear entangling gates, realize a near-deterministic cnot gate between the two nuclear spins [beukers_control_2024].
@@ III System compatibility
Figure 3: QD and G4V hardware profiles. a) Energy levels and decay paths of a quantum dot (QD). The QD includes a ground state $|0\rangle$, two approximately degenerate single-excitation states, $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$, and a biexcitation state, $|\rm XX\rangle$, which together give rise to two cascade decay paths. The decay times $T_{\rm X}$ and $T_{\rm XX}$ of the single- and biexcitation states correspond to linewidths $\gamma_{\rm X}$ and $\gamma_{\rm XX}$, respectively. Photons are emitted with horizontal (H) or vertical (V) polarization, depending on the decay path; the presence or absence of a prime symbol for the labels H and V denotes different photon frequencies. b) Frequency profile of photons emitted by the QD. According to the experiment presented in [schimpf_quantum_2021], the photons from both $|\rm XX\rangle$, $|\rm X\rangle$ and $|\rm X^{\prime}\rangle$ exhibit Lorentzian spectral profiles centered at $\omega_{\rm 0,X}$ and $\omega_{\rm 0,XX}$, with bandwidths $\gamma_{\rm XX}=8.33$ GHz and $\gamma_{\rm X}=4.34$ GHz, respectively. c) Fiber–cavity interface and integration of a group-IV-vacancy color center (G4V) into the sawfish nanophotonic crystal cavity [bopp_sawfish_2024]. The incoming photon from the fiber is in a superposition of early and late time-bin states, encoding a qubit. The cavity supports a single optical mode with resonance frequency $\omega_{c}$, coupling strength $\bm{g}=[g_{1A},g_{2B},g_{2A},g_{1B}]$ to the G4V, and a total loss rate $\kappa$. d) Energy levels of the G4V center (SiV or SnV), modeled as a four-level system with ground states $|1\rangle$ and $|2\rangle$ and excited states $|A\rangle$ and $|B\rangle$. The optical transitions $|1\rangle\leftrightarrow|A\rangle$ and $|2\rangle\leftrightarrow|B\rangle$ have linewidths $\gamma_{1A}$ and $\gamma_{2B}$, respectively. The cavity resonance $\omega_{c}$ is detuned by $\delta$ from the $|1\rangle\leftrightarrow|A\rangle$ transition. The ground-states $|1\rangle$ and $|2\rangle$ circled by the orange line make up the qubit of the G4V spin, whose splitting is denoted as $\omega_{s}$. e) Operational sequence for implementing the spin–photon entanglement gate via the reflection-based scheme (see App. A). The protocol consists of four steps: (1) initialize the G4V spin in state $|1\rangle$; (2) scatter the early time-bin photon; (3) apply a $\pi/2$ rotation around the $y$-axis to the spin; (4) scatter the late time-bin photon. Fig. e) is adapted from [bhaskar_experimental_2020].
@@
Compatibility between the QD photon emitter and the G4V single-spin quantum memory is a central challenge in our protocol. While the central frequencies of the two systems can be aligned using frequency converters, matching their bandwidths and lineshapes remains a formidable task. The QD emitter has a natural lifetime of 125125250250 ps [schimpf_quantum_2021], corresponding to a bandwidth on the order of 10 GHz. In contrast, the G4V spin has a natural lifetime of 1.71.74.54.5 ns [bopp_sawfish_2024], implying a bandwidth on the order of 100 MHz—only about 1% of that of the QD. This mismatch implies that the G4V would fail to interact efficiently with approximately 99% of the QD-emitted photons. Nevertheless, this issue can be mitigated by coupling the G4V and the broadband photon via an intermediate cavity with an optimized total energy decay rate and central mode frequency relative to the G4V transition [omlor_entanglement_2024].
Following a decay cascade from the biexcitation state to the ground state, the experimentally investgated QD considered here emits two photons with Lorentzian spectral profiles [tran_nanodiamonds_2017] and bandwidths $\gamma_{\rm XX}=8.33$ GHz and $\gamma_{\rm X}=4.34$ GHz, respectively [schimpf_quantum_2021], as shown in Fig. 3(b). Each photon can optionally pass through a Fabry–Perot interferometer for frequency filtering [Ismail2016] (see App. D.4 for details) and is routed to one of the two adjacent nodes containing a G4V spin, thereby distributing entanglement. Since the two photons have different bandwidths, using two distinct cavities optimized for each photon would yield the best performance. However, for simplicity and practical implementation, we adopt a symmetric design for both cavities, balancing system complexity and performance. As a result, only three parameters require optimization: the photon central frequency $\omega_{0}$ (as targeted by the frequency converters), the cavity resonance frequency $\omega_{c}$, and the cavity loss rate $\kappa$.
Because the challenge of low success probability of entanglement transfer, mainly caused by photon loss during fiber transmission and local optical circuits, is addressed through multiplexing, quantum state and gate fidelities become the dominant factor limiting repeater performance. Therefore, our objective is to maximize the fidelity of the resulting G4V spin-spin state relative to the raw G4V electron-electron Bell-state, $|\mathrm{Bell}\rangle=\frac{1}{\sqrt{2}}(|11\rangle+|22\rangle)$, where $|ij\rangle$ ($i,j\in\{1,2\}$) is the quantum state where the left(right) G4V electron in an elementary link is in the $|i(j)\rangle$ state. This leads to the objective function [Liang2019]
@@
To examine the quantum state transfer from one of the two photons in the pair to the G4V spin, we consider the setup shown in Fig. 3(c). Via the polarization to time-bin converter, the photonic qubit is encoded in early and late time bins. The photon is guided through a fiber into a single-sided, overcoupled nanophotonic crystal cavity [bopp_sawfish_2024], where it interacts with the G4V spin. After interaction, the photon is reflected back into the fiber and directed to a measurement device. In the simplest case spin-photon entanglement succeeds when the phase contrast fulfills $\Delta\phi(\omega):=\phi_{1}(\omega)-\phi_{2}(\omega)=\pi$ on a frequency range around the central frequency of the incoming photon where the expressions $\phi_{i}$ are the phases of the reflected photon when the spin is initialized in $|i\rangle$ for $i=1,2$. The phase contrast $\Delta\phi$ is determined by the optical splitting $\Delta\omega_{s}$ of the G4V, which corresponds to the energy difference between the optical transitions connecting the excited states $|A\rangle$ and $|B\rangle$, and the ground states $|1\rangle$ and $|2\rangle$.